OpenStudy (anonymous):

Question and Answer Down Below (Please Allow time to be written) Question type: Geometric Sequences

OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence? 6,250; 1,250; 250; 25 … \[A: A _{N}=\frac{ 1 }{ 5 }(6250)^{N-1}\] \[B: A_{N}=6250\left(\begin{matrix}1 \\ 5\end{matrix}\right)^{N-1}\] \[C: A _{N}=6250(\frac{ 1 }{ 5 })^{N-1}\] \[D: A _{N}=6250(5)^{N-1}\]